OpenStudy (anonymous):

Solve the equation 4(x – 2) – x = 3x – 8

OpenStudy (amorfide):

first, multiply out of the brackets 4(x-2)=4x-8 re write 4x-8-x=3x-8 move all x's onto one side, and all constants onto the other side, finish it from there

OpenStudy (anonymous):

Thats confusing lol

OpenStudy (anonymous):

4(x-2)-x=3x-8 4x-8-x=3x-8 4x-x-8=3x-8 3x-8=3x-8 +8 +8 ______________ 3x=3x-0 -3x -3x ___________ 0=0

OpenStudy (anonymous):

No solution? or 0?

OpenStudy (anonymous):

real number

OpenStudy (anonymous):

it is 0 because 0=0 is true, it is only no solution if it is incorect, example, 4=0 thats no solution

OpenStudy (anonymous):

What is the graph of y = x + 2?

OpenStudy (anonymous):

3x-8=3x-8 for real number example -3x-8=3x-8 +8 +8 _______________ -3x=3x-0 -3x -3x _____________ -6x=0 no solution